\begin{thebibliography}{1}
% \bibitem{AA}  N. Aydin, T. Asamov,  The $\Bbb Z_4$ Database [Online]. Available: http://z4codes.info/. accessed on 28.08.2015


\bibitem{AK}  E. F. Assmus, Jr., and J. D. Key,  ``Affine and projective planes,'' \textit{Discrete
Math.,} vol. 83, pp. 161-187, Aug. 1990.


\bibitem{BPR} P. Beelen, S. Puchinger, and J. R. n\'e Nielsen,
``Twisted Reed-Solomon codes,"
in \textit{IEEE Int. Symp. on Information Theory (ISIT),}
pp. 336-340, Jun. 2017. %OK

\bibitem{BCCGM} J. Bringer, C. Carlet, H. Chabanne, S. Guilley, H. Maghrebi, “Orthogonal direct sum masking: A smartcard friendly computation paradigm in a code, with builtin protection against side-channel and fault attacks,” in Information Security Theory and Practice. Securing the Internet of Things (Lecture Notes in Computer Science), vol. 8501. Heraklion, Greece: Springer, 2014, pp. 40-56.

\bibitem{CG} C. Carlet, S. Guilley, “Complementary dual codes for counter- measures to side-channel attacks,” in Coding Theory and Applications (CIM Series in Mathematical Sciences), vol. 3, E. R. Pinto, Ed. Cham, Switzerland: Springer-Verlag, 2014, pp. 97-105.



\bibitem{CLM} C. Carlet, C. Li,  S. Mesnager,  ``Linear codes with small hulls in semi-primitive case,'' \textit{Des. Codes Cryptogr.,} vol. 87  no. 12, pp. 3063-3075, 2019.

\bibitem{CMTQP} C. Carlet, S. Mesnager, C. Tang, Y. Qi,  R. Pellikaan, “Linear codes over $\Bbb F_q$ are equivalent to LCD codes for $q > 3$,” \textit{IEEE Trans. Inf. Theory,} vol. 64, no. 4, pp. 3010-3017, Apr. 2018.




\bibitem{CL} B. Chen and H. Liu, “New constructions of MDS codes with complementary duals,” \textit{IEEE Trans. Inf. Theory,} vol. 64, no. 8, pp. 5776-5782, Aug. 2018.

\bibitem{FFLZ} W. Fang, F.-W. Fu, L. Li, and S. Zhu, ``Euclidean and Hermitian hulls of MDS codes and their applications to EAQECCs,'' \textit{IEEE Trans. Inf. Theory,} vol. 66, no. 6, pp. 3527-3537, Jun. 2020. 

\bibitem{J} L. Jin, “Construction of MDS codes with complementary duals,” \textit{IEEE Trans. Inf. Theory,} vol. 63, no. 5, pp. 2843-2847, May 2017.



\bibitem{L1}  J. Leon, ``Computing automorphism groups of error-correcting codes,''
\textit{IEEE Trans. Inf. Theory,} vol. IT-28, no. 3, pp. 496-511, May 1982.

\bibitem{L2} J. S. Leon, ``Permutation group algorithms based on partitions, I: Theory
and algorithms,'' \textit{J. Symbolic Comput.,} vol. 12, pp. 533-583, Oct. 1991.


\bibitem{L} C. Li, “Hermitian LCD codes from cyclic codes,” \textit{Designs, Codes Cryptography,} vol. 86, no. 10, pp. 2261-2278, Oct. 2018.

\bibitem{LDL} C. Li, C. Ding, and S. Li, “LCD cyclic codes over finite fields,” \textit{IEEE Trans. Inf. Theory,} vol. 63, no. 7, pp. 4344-4356, Jul. 2017.



\bibitem{LZ} C. Li,  P. Zeng,  ``Constructions of linear codes with one-dimensional hull,'' \textit{IEEE Trans. Inf. Theory,} vol. 65, no. 3, pp.  1668-1676, Mar. 2019.


\bibitem{LLDL} S. Li, C. Li, C. Ding, and H. Liu, ``Two families of LCD BCH codes,''
\textit{IEEE Trans. Inf. Theory,} vol. 63, no. 9, pp. 5699-5717, Sep. 2017.


\bibitem{LCC} G. Luo, X. Cao, and X. Chen, ``MDS codes with hulls of arbitrary dimensions and their quantum error correction,” \textit{IEEE Trans. Inf. Theory,} vol. 65, no. 5, pp. 2944-2952, May 2019.



\bibitem{MS} F. J. MacWilliams and N. J. A. Sloane,
\textit{The Theory of Error-Correcting Codes,}  Amsterdam, The Netherlands: North Holland, 1977. %OK

\bibitem{M} J. L. Massey, “Linear codes with complementary duals,” \textit{Discrete Math.,} 106-107: pp. 337–342, Sep. 1992.


\bibitem{RS} I. S. Reed and G. Solomon,
``Polynomial codes over certain finite fields,"
\textit{J. Soc. Ind. Appl. Math.,}
vol. 8, no. 2, pp. 300-304, Jun. 1960. %OK

%\bibitem{R}  R. M. Roth, Introduction to Coding Theory. CUP, 2006.

\bibitem{RL} R. M. Roth and A. Lempel,
``A construction of non-Reed-Solomon type MDS codes,"
\textit{IEEE Trans. Inf. Theory,}
vol. 35, no. 3, pp. 655-657, May 1989. %OK

\bibitem{RS2}  R. M. Roth and G. Seroussi,
``On generator matrices of MDS codes (Corresp.),"
\textit{IEEE Trans. Inf. Theory,}
vol. 31, no. 6, pp. 826-830, Nov. 1985. %OK
















\bibitem{S1}  N. Sendrier, ``On the dimension of the hull,'' \textit{SIAM J. Discrete Math.,} vol. 10, no. 2, pp. 282-293, 1997.



\bibitem{S2} N. Sendrier, ``Finding the permutation between equivalent linear codes:
The support splitting algorithm,'' \textit{IEEE Trans. Inf. Theory,} vol. 46, no. 4, pp. 1193-1203, Jul. 2000.

\bibitem{SS} N. Sendrier and G. Skersys, ``On the computation of the automorphism
group of a linear code,'' in \textit{Proc. IEEE Int. Symp. Inf. Theory,}
Washington, DC, USA, p. 13,  Jun. 2001.


\bibitem{WHL}  Y. Wu, J. Y. Hyun, Y. Lee, ``New LCD MDS codes of non-Reed-Solomon type,'' submitted for publication.



\bibitem{WY}  Y. Wu, Q. Yue, ``Factorizations of binomial polynomials and enumerations of LCD and self-dual constacyclic codes,''  \textit{IEEE Trans. Inf. Theory,} vol. 65, no.3, pp. 1740-1751, Mar. 2019.


\bibitem{WYF1}  Y. Wu, Q. Yue, S. Fan, ``Further factorization of $x^n -1$ over a finite field,'' \textit{Finite Fields Appl.,}
vol. 54, pp. 197-215, Nov. 2018.

\bibitem{WYF2} Y. Wu, Q. Yue, S. Fan, ``Self-reciprocal and self-conjugate-reciprocal irreducible factors of $x^n -\lambda$  and their applications,'' \textit{Finite Fields Appl.,} vol. 63, 101648, Mar. 2020.





\bibitem{WYZY}  Y. Wu, Q. Yue, X. Zhu, S. Yang, ``Weight enumerators of reducible cyclic codes and their dual codes,''
\textit{Discrete Math.,} vol. 342, no.3, pp. 671-682, Mar. 2019.

%\bibitem{CH} S. Chang,  J.Y. Hyun,  Linear codes from simplicial complexes,  Des. Codes Cryptogr.,  86:  2167-2181, 2018.

 %  \bibitem{CDY}  C. Carlet, C. Ding, J. Yuan, Linear codes from perfect nonlinear mappings and their secret sharing schemes, IEEE Trans. Inf. Theory, 51: 2089-2102  (2005)

% \bibitem{D1} C. Ding, Linear codes from some 2-designs, IEEE Trans. Inf. Theory, 61(6):  3265-3275, 2015.

%  \bibitem{DHKW} C. Ding, T. Helleseth, T. Kl\o ve, X. Wang, A general construction of  Cartesian authentication codes, IEEE Trans. Inf. Theory, 53: 2229-2235, 2007.

% \bibitem{DN} C. Ding, H. Niederreiter, Cyclotomic linear codes of order 3, IEEE Trans. Inf. Theory, 53(6):  2274-2277, 2007.

%\bibitem{G2} M. Grassl,  Bounds on the minimum distance of linear codes.  http://www.codetables.de.


% \bibitem{G2} Grassl M.:  Bounds on the minimum distance of linear codes.  http://www.codetables.de.


 %\bibitem{G} J. H. Griesmer, A bound for error correcting codes, IBM J. Res. Dev., 4: 532–542, 1960.

% \bibitem{HY1}  Z. Heng, Q. Yue,  A class of binary linear codes with at most three weights,   IEEE Commun. Lett., 19: 1488-1491,  2015.


 %  \bibitem{HY2}  Z. Heng, Q. Yue, Two classes of two-weight linear codes, Finite Fields Appl.,  38: 72-92, 2016.

%  \bibitem{HY3}  Z. Heng,  Q. Yue, Evaluation of the Hamming weights of a class of linear codes based on Gauss sums,  Des. Codes Cryptogr., 83: 307-326, 2017.



%  \bibitem{HYL} Z. Heng, Q. Yue, C. Li,  Three classes of linear codes with two or three weights, Discrete Math., 339:  2832-2847, 2016.

%\bibitem{H} A. R. Hammons Jr., P. V. Kumar, A. R. Calderbank, N. J. A. Solance, P. Sol$\acute{e}$, The $\Bbb Z_4$ linearity of
%Kerdock, Preparata, Goethals and related codes, IEEE Trans. Inf. Theory, 40:   301–319, 1994.

%    \bibitem{HP}  W. C. Huffman, V. Pless, Fundamentals of Error-Correcting Codes, Cambridge University Press, Cambridge, 2003.

    %  \bibitem{HLL} J. Y. Hyun,  J. Lee,  Y. Lee, Infinite families of optimal linear codes constructed from simplicial complexes, submitted for publication.

 %    \bibitem{HKWY} J. Y. Hyun,  H. K. Kim, Y. Wu, Q. Yue, Optimal minimal linear codes from posets, submitted for publication.

 %\bibitem{N}  J. Neggers, H. S. Kim,  Basic posets, World Scientific, 1998.

   %\bibitem{HLL} J. Y. Hyun,  J. Lee, Y. Lee, Infinite families of optimal linear codes constructed from simplicial complexes, submitted for publication.




%  \bibitem{LYL}  C. Li, Q. Yue,   F. Li,  Weight distributions of cyclic codes with respect to pairwise coprime order elements, Finite Fields Appl.,  28: 94-114, 2014.

%  \bibitem{LM}  H. Liu,  Y. Maouche, Two or few-weight trace codes over $\Bbb F_q+u\Bbb F_q$,  IEEE Trans. Inf. Theory, 65(5):  2696- 2703, 2019.

%  \bibitem{LM2}  H. Liu,  Y. Maouche, Several new classes of linear codes with few weights, Cryptogr. Commun., 11: 137-146, 2019.



%   \bibitem{LL} Y. Liu, Z. Liu, On some classes of codes with a few weights,  Adv. Math. Commun., 12: 415-428, 2018.




% \bibitem{AB} Ashikhmin A.,  Barg A.:  Minimal vectors in linear codes. IEEE Trans. Inf. Theory  {\bf 44}, 2010-2017 (1998)

 %\bibitem{BB} Bartoli D., Bonini M.:  Minimal linear codes in odd characteristic.  IEEE Trans. Inf. Theory  DOI: 10.1109/TIT.2019.2891992 (2019)

%\bibitem {CD} C. Carlet and C. Ding, Nonlinearities of S-boxes, Finite Fields Appl, 13: 121-135, 2007.



%\bibitem{CMP} Cohen G., Mesnager S., Patey A.: On minimal and quasi-minimal linear codes, in: Proceedings of IMACC (Lecture Notes in Computer Science, vol. 8308), M. Stam, Eds. Berlin: Springer-Verlag, 85-98 (2003)

%\bibitem{D1} C. Ding, Linear codes from some 2-designs, IEEE Trans. Inform. Theory, 61(6):  3265–3275, 2015.

%\bibitem{D2}  Ding C.:  A construction of binary linear codes from Boolean functions. Discrete Math. {\bf 339}, 2288-2303 (2016)

%\bibitem{DD}  K. Ding, C. Ding,  Binary linear codes with three weights, IEEE Commun. Lett., 18(11):1879-1882, 2014.

%\bibitem{DHZ}  Ding C., Heng Z., Zhou Z.: Minimal binary linear codes.  IEEE Trans. Inf. Theory  {\bf 64},  6536-6545 (2018)

%\bibitem{DY}  Ding C., Yuan J.: Covering and secret sharing with linear codes, in: Discrete Mathematics and Theoretical Computer Science, Lecture Notes in Computer Science 2731, Springer Verlag, 11-25 (2003)



%\bibitem{G} Griesmer J.H.:  A bound for error correcting codes. IBM J. Res. Dev. {\bf 4}, 532–542 (1960)



%\bibitem{G}   J.H. Griesmer, A bound for error-correcting codes, IBMJ. Res. Dev. 4: 532-542, 1960.

%\bibitem{HDZ} Heng Z.,  Ding C., Zhou Z.: Minimal  linear codes over finite fields.  Finite Fields Appl. {\bf 54} 176-196 (2018)





%\bibitem{HY} Z. Heng , Q. Yue,  A construction of $q$-ary linear codes with two weights, Finite Fields Appl.,  48: 20-42, 2017.

%\bibitem{LYL}  C. Li, Q. Yue,  Fengwei Li. Weight distributions of cyclic codes with respect to pairwise coprime order elements. Finite Fields Appl., 28: 94-114, 2014.





%\bibitem{M}   Massey J. L.: Minimal codewords and secret sharing, in: Proc. 6th Joint Swedish-Russian Workshop on Information Theory, M\"olle, Sweden, 246-249 (1993)

%\bibitem{M2} Mesnager S.: Linear codes with few weights from weakly regular bent functions based on a generic construction. Cryptogr. Commun. {\bf 9}, 71-84 (2017)

%\bibitem{N}  J. Neggers, H. S. Kim,  Basic posets, World Scientific, 1998.

%\bibitem{WHWK}  Wadayama T., Hada T., Wakasugi K., Kasahara M.: Upper and lower bounds on maximum nonlinearity of $n$-input $m$-output Boolean function. Des. Codes Cryptogr.  {\bf 23},   23-33 (2001)

%\bibitem{WDX}   Q. Wang, K. Ding, R. Xue,  Binary linear codes with two weights, IEEE Commun. Lett.,  19(7): 1097-1100, 2015.

%\bibitem{YD} Yuan J., Ding C.: Secret sharing schemes from three classes of linear codes. IEEE Trans. Inf. Theory {\bf 52}, 206-212 (2006)

%\bibitem{YKT} S. Yang, X. Kong, C. Tang, A construction of linear codes and their
%complete weight enumerators, Finite Fields Appl., 48: 196-226, 2017.

%\bibitem{YY}  S. Yang, Z.-A. Yao, Complete weight enumerators of a class of linear codes, Discrete Math.,  340: 729-739,   2017.

%\bibitem{ZTLD}  Z. Zhou, C. Tang, X.  Li, C. Ding,  Binary LCD codes and self-orthogonal codes from a generic construction, IEEE Trans. Inf. Theory, 65(1): 16-27, 2019.
%\bibitem{B} F. E. Brochero Mart\'{\i}nez,  Structure of finite dihedral group algebra, Finite Fields Appl., 2015, 35: 204-214.



%\bibitem{BA} S. Batra, S.  Arora, Some cyclic codes of length $2p^n$, Designs, Codes and Cryptography, 2011, 61: 41-69.

% \bibitem{CLZ} B. Chen, H. Liu, G. Zhang,  A class of minimal cyclic codes over finite fields, Designs, Codes and Cryptography, 2013, doi:10.1007/s10623-013-9857-9.


% \bibitem{C16} Chen B, Ling S, Zhang G. Enumeration formulas for self-dual cyclic codes[J]. Finite Fields and Their Applications, 2016, 42: 1-22.

% \bibitem{F00} Feng, K. Algebraic Number Theory[M]. Science Press, 2000.


% \bibitem{FM} R. Ferraz, C. Milies,  Idempotents in group algebras and minimal abelian codes, Finite Fields Appl.,  2007, 13: 382-393.


%\bibitem{GCL} Y. Gao, J. Chen, Y. Li,  Some $\ast$-clean group rings,  Algebra Colloq., 2015, 22 (1): 169-180.


%  \bibitem{HKCS94} Hammons  A, Kumar P , Calderbank A , Sloane N, and Sol\'e P. The  $\mathbb{Z}_4$-linearity of Kerdock, Preparata, Goethals, and Related codes[J]. Transactions on Information Theory, 1994, 40: 301-319.


 %\bibitem{HRZ}D. Han, Y. Ren, H. Zhang,  On $\ast$-clean group rings over abelian groups,  J. Algebra Appl., 2016: 1750152.


% \bibitem{HLY}  H. Huang, Y. Li, P. Yuan,  On $\ast$-clean group rings II, Comm. Algebra, 2016, 44(7): 3171-3181.

 % \bibitem{HLT} H. Huang, Y. Li, G. Tang,  On $\ast$-clean non-commutative group rings, J. Algebra Appl., 2016, 15(08): 1650150.

%\bibitem{I}  N. Immormino, Clean rings and clean group rings, Ph.D. Dissertation, Bowling Green State University, 2013.

%\bibitem{IM14} N. Immormino,  W. McGovern, Examples of clean commutative group rings, J. Algebra, 2014, 405: 168-178.



%\bibitem{KLP} P. Kanwar, S. L\'opez-Permouth,  Cyclic codes over the integers modulo $p^ m$, Finite Fields Appl., 1997, 3(4): 334-352.

% \bibitem{LZ} C. Li, Y. Zhou, On strongly $\ast$-clean rings, J. Algebra Appl., 2011, 10 (6): 1363-1370.


% \bibitem{12} Li F. Primitive idempotents in group algebras and minimal abelian codes[J]. Journal of Applied Mathematics and Computing, 2015, 49(1-2): 1-11.

% \bibitem{LYL1} F.  Li, Q. Yue, C. Li,  The minimum  Hamming distances of  irreducible cyclic codes,  Finite Fields Appl., 2014, 29:  225-242.

% \bibitem{LYL2}  F.  Li, Q. Yue, C. Li, Irreducible cyclic codes of length $4p^n$ and $8p^n$, Finite Fields Appl., 2015, 34: 208-234.

%\bibitem{LY} F.  Li, Q. Yue,  The primitive idempotents and weight distributions of  irreducible constacyclic codes, Des. Codes Cryptogr., 2018, 86: 771-784.

%\bibitem{LPY} Y.  Li, M.  Parmenter,  P. Yuan.   On $\ast$-clean group rings, J. Algebra Appl., 2015, 14(1): 1550004.

%\bibitem{L} R. Lidl , Niederreiter H.  Finite fields, Cambridge University press, 1997.

%\bibitem{L}  Van Lint J H. Introduction to coding theory[M]. Springer Science $\&$ Business Media, 2012.

%\bibitem{LYW} Y. Lin, Q. Yue,  Y. Wu, Primitive idempotents of irreducible cyclic codes of length, Math. Probl. Eng., 2018, Article ID 6962508, 9 pages, 2018. https://doi.org/10.1155/2018/6962508.


%\bibitem{N}  W. Nicholson. Lifting idempotents and exchange rings, Trans. Amer. Math. Soc., 1977, 229: 269-278.



% \bibitem{PA} M. Pruthi, S. Arora,  Minimal codes of prime power length, Finite Fields Appl., 1997, 3: 99-113.

 %\bibitem{PA} M. Pruthi, S. K. Arora,  Minimal cyclic codes of prime power length, Finite Fields  and Their Applications, 1997, 3:  99-113.
%\bibitem{P}  Pruthi M. Cyclic codes of length $2^m$[J], Proc. Indian Acad. Sci. Math. Sci, 2001,  111: 371-379.



%\bibitem{QK14} Qian K, Zhu S, Kai X. On cyclic self-orthogonal codes over $\mathbb{Z}_{2m}$ , Finite Fields Appl. 2014, 33 :
%53-65.
% \bibitem{S3} M. Shi, Y. Guan,  P. Sol$\acute{e}$, Two new families of two-weight codes,
%IEEE Trans. Inf. Theory, 63(10):  6240–6246, 2017.

%\bibitem{S1} M. Shi, Y. Liu, P. Sol$\acute{e}$, Optimal two-weight codes from trace codes
%over $\Bbb F_2+ u\Bbb F_2$,  IEEE Commun. Lett., 20(12): 2346-2349, 2016.

%\bibitem{S2}  M. Shi, R. Wu, Y. Liu,  P. Sol$\acute{e}$, Two and three weight codes over
%$\Bbb F_2+ u\Bbb F_2$, Cryptogr. Commun., 9(5):  637-646, 2017.




%\bibitem{W} Z. X. Wan, Quaternary Codes, World Scientific, Singapore, 1997.

%\bibitem{WHY} Y. Wu, J. Y.  Hyun, Q. Yue,   Binary optimal linear codes from posets of the
%disjoint union of two chains, arXiv:1906.12017.

%\bibitem{WZY} Y. Wu, X. Zhu, Q. Yue,   Optimal few-weight codes from simplicial complexes, submitted for publication.








%PTIMAL FEW-WEIGHT CODES FROM SIMPLICIAL
%COMPLEXES
%\bibitem{SP} Singh R, Pruthi M. Primitive idempotents of irreducible quadratic residue cyclic codes of length $p^nq^m$[J], Int. J. Algebra, 2011, 5:  285-294.

%\bibitem{TWL} G. Tang, Y. Wu, Y. Li,  $\ast$-Cleanness of finite group rings, Comm. Algebra, 2017, 45(10): 4190-4195.



%\bibitem{V} L.  Va\u{s} ,  $\ast$-Clean rings; some clean and almost clean Baer $\ast$-rings and Von Neuman
%algebra, J. Algebra, 2010, 324:  3388-3400

%\bibitem{W} Z. Wan,  Lectures on finite fields and Galois rings[M]. World Scientific, 2003.


% \bibitem{WYL} Y. Wu, Q. Yue, F. Li,  Primitive idempotents of irreducible cyclic codes and self-dual cyclic codes over Galois rings, Discrete Math., 2018, 341(6): 1755-1767.

%  \bibitem{YKT} S. Yang, X. Kong, C. Tang, A construction of linear codes and their complete weight enumerators, Finite Fields Appl., 2017, 48: 196–226.


% \bibitem{YY}  S. Yang, Z.-A. Yao, Complete weight enumerators of a class of linear codes, Discrete Math.,
%2017, 340: 729-739.


 %\bibitem{YYZ}   S. Yang, Z.-A. Yao, C.-A. Zhao, The weight distributions of two classes of p-ary cyclic codes
%with few weights, Finite Fields Appl., 2017, 44: 76-91.

%\bibitem{13}Berman,S.D. :Semiple cyclic and abelian codes.II. Kybernetika3,21-30(1967)






%\bibitem{14} Macwilliams, F.J. :Binary codes which are ideals in the group algebra of an abelian group. BellSyst.Tech. J. 44, 987-1011 (1970).





%\bibitem{19 }Chabanne, H.:Permutation decoding of abelian codes.IEEETrans.Inf.Theory38,1826-1829(1992).



%\bibitem{21 } Jitman,S., Ling,S., Liu,H., Xie,X.: Abelian codes in principal ideal group algebra. IEEETrans. Inf.Theory 59, 3046-3058 (2013).





%\bibitem{CH} S. Chang, J. Y. Hyun, Linear codes from simplicial complexes, Des. Codes Cryptogr., 86: 2167-2181, 2018.


 % \bibitem{D2}  C. Ding, A construction of binary linear codes from Boolean functions, Discrete Math., 339, 2288-
%2303, 2016.

% \bibitem{DD}  K. Ding, C. Ding,  Binary linear codes with three weights, IEEE Commun. Lett., 2014, 18(11):1879-1882.











   % \bibitem{HP}  W. C. Huffman, V. Pless, Fundamentals of Error-Correcting Codes, Cambridge University Press, Cambridge, 2003.







 %    \bibitem{N}  J. Neggers, H. S. Kim,  Basic posets, World Scientific, 1998.

   %\bibitem{WDX}   Q. Wang, K. Ding, R. Xue,  Binary linear codes with two weights, IEEE Commun. Lett.,  19(7), 1097-1100 2015.


  % \bibitem{DD1}  K. Ding, C. Ding, Binary linear codes with three weights,  IEEE Commun. Lett., 18: 1879-1882, 2014.

%  \bibitem{DD2}  K. Ding, C. Ding,  A class of two-weight and three-weight codes and their applications in secret sharing,  IEEE Trans. Inf. Theory,  61: 5835-5842, 2015.



 %   \bibitem{WDLX} Q. Wang, Ding K.,  D. D. Lin, R. Xue,  A kind of three-weight linear codes, Cryptogr.
%Commun. , 9: 315-322, 2017.

% \bibitem{WDX}   Q. Wang, K. Ding, R. Xue,  Binary linear codes with two weights, IEEE Commun. Lett., 19: 1097-1100, 2015.

 %\bibitem{YKT} S. Yang, X. Kong, C. Tang, A construction of linear codes and their complete weight enumerators,  Finite Fields Appl., 48: 196-226,  2017.

%\bibitem{YZ} S. Yang, Z.-A. Yao, Complete weight enumerators of a class of linear codes,  Discrete Math.,  340: 729-739, 2017.

% \bibitem{YD} J. Yuan, C. Ding, Secret sharing schemes from three classes of linear
%codes, IEEE Trans. Inf. Theory, 52(1):  206-212 (2006)


 %  \bibitem{ZTLD}  Z. Zhou, C. Tang, X.  Li, C. Ding,  Binary LCD codes and self-orthogonal codes from a generic construction, IEEE Transactions on Information Theory, 65(1): 16-27, 2019.





%\bibitem{23 } Rajan, B.S., Siddiqi, M.U.: Transform domain characterization of abelian codes. IEEE Trans. Inf.Theory 38(6), 1817-1821 (1992).


% \bibitem{24 }Singh, R., Pruthi, M.: Primitive idempotents of irreducible quadratic residue cyclic codes of length$p^nq^m$. Int. J. Algebra 5, 285-294 (2011).





%\bibitem{AAC} T. Andreescu, D. Andrica and I. Cucurezeanu, An Introduction to Diophantine Equations: A Problem-Based Approach, Birkh\"{a}user, 2010.


%\bibitem{AC14} D. Andrica and G. C\u{a}lug\u{a}reanu, A nil-clean $2 \times 2$ matrix over the integers which is not clean, {\it J. Algebra Appl.}, {\bf 13}(6) (2014) 1450009.

%\bibitem{BCDM13} S. Breaz, G. C\u{a}lug\u{a}reanu, P. Danchev and T. Micu,  Nil-clean matrix rings, {\it Linear Alg. Appl.}  {\bf 439} (2013), 3115-3119.
%\bibitem{CYZ} J. Chen, X. Yang and Y. Zhou, On strongly clean matrix and triangular matrix rings,  {\it Comm. Algebra} {\bf 34} (2006), 3659-3674.
%\bibitem{CH10} Dragana Cvetkovic-Ilic and Robin Harte, On Jacobson’s lemma and Drazin invertibility, {\it Appl. Math. Lett.} {\bf 23}(4) (2010), 417-420.


%\bibitem{D06} A.J. Diesl, Classes of strongly clean rings, PhD Thesis, University of California, Berkeley (2006).

%\bibitem{D13} A.J. Diesl, Nil clean rings, {\it J. Algebra}, {\bf 383} (2013), 197-211.

%\bibitem{HTY88} Y. Hirano, H. Tominaga and A. Yaqub, On rings in which every element is uniquely expressible as a sum of a nilpotent element and a certain potent element, {\it Math. J. Okayama Univ.} {\bf 30} (1988), 33-40.
%\bibitem{KL04}  D. Khurana and T.Y. Lam,  Clean matrices and unit-regular matrices, {\it J. Algebra} {\bf 280} (2)(2004), 683-698.

%\bibitem{KLZ14} M.T. Kos\c{a}n, T-K. Lee and Y. Zhou, When is every matrix over a division ring a sum of an idempotent and a nilpotent?  {\it  Linear Alg. Appl.}  {\bf 450} (2014), 7-12.


%\bibitem{KWZ16} T. Ko\c{s}an, Z. Wang and Y. Zhou, Nil-clean and strongly nil-clean rings, {\it J. Pure Appl. Algebra}, {\bf 220}(2) (2016), 633-646.

%\bibitem{L}T.Y. Lam, A First Course in Noncommutative Rings, Second Edition, Grad. Texts in Math., vol. {\bf 131}, Springer-Verlag, New York, 2001.

%\bibitem{LP14} T.Y. Lam and Pace P. Nielsen, Jacobson's lemma for Drazin inverses, {\it Contemporary Mathematics}, {\bf  609}, pp. 185-195, 2014.

%\bibitem{LH} Jr. Lenstra, W. Hendrik, Solving the Pell equation, {\it Notices. AMS}, {\bf 49}(2) (2002), 182-192.


%\bibitem{M16} 	J. Matczuk, Conjugate (nil) clean rings and K\"{o}ethe's problem,  {\it J. Algebra Appl.} {\bf 16}(2) (2017) 1750073.
%\bibitem{MSS15} W.Wm. McGovern, S. Raja and A. Sharp, Commutative nil clean group rings, {\it J. Alg. Appl.} {\bf 14}(6) (2015) 1550094.

%\bibitem{Na}  I. Nagell, Introduction to Number Theory (John Wiley, New York, Stockholm, 1951).

%\bibitem{N77} W. Nicholson, Lifting idempotents and exchange rings, {\it Trans. AMS.}, {\bf 229} (1977), 269-278.

%\bibitem{STZ} S. Sahinkaya, G. Tang and Y. Zhou, Nil-clean group rings, {\it J. Algebra Appl.}, DOI: 10.1142/S0219498817501353.
%\bibitem{TZ13} G. Tang and Y. Zhou,  A class of formal matrix rings, {\it Linear Algebra Appl.}, {\bf 438}(12) (2013), 4672-4688.
%\bibitem{ZCC12} G. Zhuang, J. Chen and J. Cui, Jacobson's lemma for the generalized Drazin inverse,{\it Linear Algebra Appl. } {\bf 436} (3) (2012),  742-746.
\end{thebibliography}

\documentclass[journal]{IEEEtran}
\usepackage{mathrsfs}
\usepackage{tabu}
\usepackage{bbm}
\usepackage{amsmath,amscd,amssymb,latexsym}
\usepackage[all]{xy}
\usepackage{mathrsfs}
\usepackage{pgf,tikz}
\usepackage{tabu}
\usepackage{graphicx}
\usepackage{multicol}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{definition}[theorem]{Definition}
\newtheorem{example}[theorem]{Example}

\newtheorem{remark}[theorem]{Remark}

\newlength{\halfpagewidth}
\divide\halfpagewidth by 2

\ifCLASSINFOpdf
\else
\fi
\begin{document}
%
\title{ Twisted Reed-Solomon Codes With One-dimensional Hull }
%
%
%

\author{Yansheng Wu~
 \thanks{Manuscript received August 31, 2020; accepted September  16, 2020. This research was supported by the National Natural Science Foundation of China (No.
61772015). The associate editor coordinating the review of this letter and approving it for publication was  Yu-Chih Huang. }
\thanks{Yansheng Wu is with the  School of Computer Science, Nanjing University of Posts and Telecommunications, Nanjing, 210023, P. R. China (e-mail: wysasd@163.com).}

\thanks{Digital Object Identifier ~~~~~~~~~}
}

\maketitle

\begin{abstract} The hull of a linear code is defined to be the intersection
of the code and its dual.  When the size of the hull is small, 
it has been proved that some algorithms for checking permutation equivalence
of two linear codes and computing the automorphism
group of a linear code are very effective in general.  Maximum distance separable (MDS) codes are codes meeting the Singleton bound. Twisted Reed-Solomon codes is a  generalization of   Reed-Solomon codes, which is also a nice construction for MDS codes. In this short letter, we obtain some  twisted Reed-Solomon  MDS codes with one-dimensional hull. Moreover, these codes  are not monomially equivalent to  Reed-Solomon codes.


\end{abstract}

\begin{IEEEkeywords}
twisted Reed-Solomon codes, one-dimensional hull, monomially equivalent.
\end{IEEEkeywords}

%
\IEEEpeerreviewmaketitle

\section{Introduction}
%
%
%
%

\IEEEPARstart{G}{iven}  a linear code $\mathcal C$ of length $n$ over the finite field $\Bbb F_q$,
the  dual code of $\mathcal C$  is defined by
$$\mathcal C^{\bot} = \{ {\bf x}\in \Bbb F_q^{n} \mid {\bf x}{\bf y} ^T=0\mbox{ for  all } {\bf y}\in \mathcal C\}$$
where ${\bf x}{\bf y} ^T$ denotes the standard inner product of two vectors ${\bf x}$ and ${\bf y}$. 
The hull of the linear code $\mathcal{C}$ is defined to be
$$\mbox{Hull}(\mathcal{C}) := \mathcal{C} \cap  \mathcal{C}^{\bot}.$$
It is clear that  $\mbox{Hull}(\mathcal{C})$ is also a linear code over $\Bbb F_q$. The hull was originally introduced in 1990 by Assmus, Jr. and Key \cite{AK} to classify finite projective planes.
It had been
shown that the hull plays an important role in determining
the complexity of algorithms for checking permutation equivalence
of two linear codes and computing the automorphism
group of a linear code (see \cite{L1,L2}, \cite{S1}-\cite{SS}), which are very effective in general when the dimension of the hull is small.

It is worth mentioning that the special case of the hulls of linear codes is of much interest. Namely the codes with trivial intersection with its dual, which is also named linear complementary dual (LCD) codes. Massey \cite{M} first introduced this class of codes and proved that there exist asymptotically good LCD codes. A practical application of binary LCD codes against side-channel attacks (SCAs) and fault injection attacks (FIAs) was investigated by Carlet et al. \cite{BCCGM} and Carlet and Guilley \cite{CG}. Since then, the study of LCD codes is thus becoming a hot research topic in coding theory (\cite{CMTQP}-\cite{J}, \cite{L,LDL,LLDL,LCC},\cite{WY}-\cite{WYZY}).
Some nice progress on linear codes with small hulls has been made, for examples (\cite{CLM, LZ}).

A {\em maximum distance separable} (MDS) code has the greatest error correcting capability
when its length and dimension is fixed. MDS codes are extensively used in communications (for example,
 Reed-Solomon codes are all MDS codes), and they have good applications in minimum storage codes and quantum codes.  There are many known constructions for MDS codes; for instance, {\em Generalized Reed-Solomon} (GRS) codes \cite{RS}, based on the equivalent problem of finding $n$-arcs in projective geometry \cite{MS}, circulant matrices \cite{RL}, Hankel matrices \cite{RS2}, or extending GRS codes.

Recently the authors in \cite{FFLZ,LCC}  investigated the hull of MDS codes via generalized Reed-Solomon codes over finite fields. Beelen {\em et al.} \cite{BPR} first gave the definition of  twisted Reed-Solomon codes, which is a generalization of the Reed-Solomon codes, and they proved under some conditions twisted Reed-Solomon codes could be not monomially equivalent to the Reed-Solomon codes.  However, the hull of twisted Reed-Solomon codes have not been studied in that paper. Recently, Wu, Hyun and Lee \cite{WHL} constructed some LCD twisted Reed-Solomon codes.

 
  In this letter, as a follow-up work we will focus on the hull of twisted Reed-Solomon codes. In particular, we will consider to construct some  twisted Reed-Solomon MDS codes with one-dimensional hull, which are not monomially equivalent to Reed-Solomon codes. 
The rest of this letter is organized as follows. In Section II, we introduce basic concepts on the hull of linear codes and twisted Reed-Solomon codes. In Sections III, we present our main results and give some examples. 
We  conclude the letter in Section IV.

\section{Preliminaries}
Let $\Bbb F_q$ be the finite field of order $q$, where $q$ is a prime power.
An $[n, k]_q$ linear code $\mathcal{C}$ over $\Bbb F_q$  is a $k$-dimensional
subspace of $\Bbb F_q^n$. The minimum distance d of
a linear code $\mathcal{C}$ is bounded by the so-called {\em Singleton bound }: $d\le n-k+1$.
If $d= n-k+1$, then the code $\mathcal{C}$ is called a {\em maximum distance separable} (MDS) code.

The following lemma on the hull of linear codes, which is very important for obtaining our main results.

\begin{lemma}\cite[Proposition 1]{LZ} Let $\mathcal C$ be an $[n, k]$ linear code over $\Bbb F_q$ with
generator matrix $G$. Then the code $\mathcal C$ has one-dimensional hull if and only if the rank of the matrix $GG^T$ is  $k-1$, where $G^T$ denotes the transpose of $G$.
\end{lemma}

Recall that a \emph{monomial matrix} is a square matrix which has exactly one nonzero entry in
each row and each column. 

\begin{definition}{\rm Let $\mathcal C_1$ and $\mathcal C_2$ be two linear codes  of the same length over $\Bbb F_q$,  and let $G_1$ be a generator matrix of $\mathcal C_1$. Then
 $\mathcal C_1$ and $\mathcal C_2$ are {\em monomially equivalent} if
 there is a monomial matrix $M$ such that $G_1M$ is a generator matrix of $\mathcal C_2$.
}
 \end{definition}

Next we will recall some constructions of MDS codes.
We begin with the well-known generalized Reed-Solomon codes.

 \begin{definition}{\rm \label{def1} Let $\alpha_{1},\ldots,\alpha_{n}$ be distinct elements in $\Bbb F_{{q}} \cup \{\infty\}$  and $v_{1},\ldots,v_{n}$ be nonzero elements in $\Bbb F_{{q}}$. For $1\leq k \leq n$, the corresponding {\em generalized Reed-Solomon $(GRS)$ code} over $\Bbb F_{{q}}$ is defined by
$$\resizebox{9cm}{!}{ $GR{S_k}({\boldsymbol{ \alpha}},{\bf v}): = \left\{({v_1}f({\alpha_1}), \ldots ,{v_{n }}f({\alpha_{n }})) \mid f(x) \in {\Bbb F_{q}}[x],\phantom{.} \deg(f(x)) < k \right\},$}$$ where $\boldsymbol{ \alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})\in (\Bbb F_{{q}}\cup \{\infty\})^{n}$ and ${\bf v}=(v_{1},v_{2},\ldots,v_{n})$, and the quantity $f(\infty)$ is defined as the coefficient of $x^{k-1}$ in the polynomial $f$. 
}
 \end{definition}

If $v_i=1$ for every $i=1, \ldots, n$, then $GR{S_k}({\boldsymbol{ \alpha}},{\bf v})$ is called a {\it Reed-Solomon code.} In fact,   $GRS_k(\boldsymbol{\alpha},{\bf v})$ has a generator matrix  as follows: $$\resizebox{9cm}{!}{ $\left(\begin{array}{cclc} v_1 &v_2 &\ldots  &v_n\\ v_1\alpha_1 &v_2\alpha_2 &\ldots & v_n\alpha_{n}\\ \vdots &\vdots &\ddots&\vdots \\ v_1\alpha_1^{k-1} &v_2\alpha_2^{k-1} &\ldots & v_n\alpha^{k-1}_{n}\end{array}\right)=\left(\begin{array}{cclc} 1 &1 &\ldots  &1\\ \alpha_1 &\alpha_2 &\ldots &\alpha_{n}\\ \vdots &\vdots &\ddots&\vdots \\ \alpha_1^{k-1} &\alpha_2^{k-1} &\ldots & \alpha^{k-1}_{n}\end{array}\right) \left(\begin{array}{cclc} v_1 &0 &\ldots  &0\\ 0 &v_2 &\ldots &0\\ \vdots &\vdots &\ddots&\vdots \\ 0 &0 &\ldots & v_{n}\end{array}\right).$}$$
 It is well-known that a generalized Reed-Solomon code $GRS_k(\boldsymbol{\alpha},{\bf v})$ is an  $ {\left[ {n,k,n-k+1} \right] }$ MDS code and it is monomially equivalent to a Reed-Solomon code.

In 2007, Beelen {\em et al.} \cite{BPR} presented a generalization of  Reed-Solomon codes, so-called {\it twisted Reed-Solomon codes}.

 \begin{definition} \label{def2} Let $\eta$ be a nonzero element in the finite field $\Bbb F_q$. Let $k,t$ and $h$ be nonnegative integers such that $0\le h<k\le q$, $k<n$, and $0<t\le n-k$.  Let $\alpha_{1},\ldots,\alpha_{n}$  be distinct elements in $\Bbb F_{{q}} \cup \{\infty\}$, and we write $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})$. Then the corresponding {\em twisted Reed-Solomon code} over $\Bbb F_q$ of length $n$ and dimension $k$ is given by
\begin{eqnarray*}\resizebox{9cm}{!}{ $\mathcal C_k(\boldsymbol{\alpha}, t,h,\eta)=
\{(f(\alpha_1),  \cdots, f(\alpha_{n})): f(x)=\sum_{i=0}^{k-1}a_ix^i+\eta a_hx^{k-1+t}\in \Bbb F_q[x]\}$}.\end{eqnarray*}

 \end{definition}

 In fact,  \begin{equation}\resizebox{8cm}{!}{ $G=\left(\begin{array}{cclcc}
1 &1 &\ldots  &1\\
 \alpha_1 &\alpha_2 &\ldots & \alpha_{n}\\
   \vdots &\vdots &\ddots&\vdots \\
   \alpha_1^{h-1} &\alpha_2^{h-1} &\ldots & \alpha^{h-1}_{n}\\
   \alpha_1^h+\eta \alpha_1^{k-1+t} &\alpha_2^h+\eta \alpha_2^{k-1+t} &\ldots & \alpha_{n}^h+\eta \alpha_n^{k-1+t}\\
  \alpha_1^{h+1} &\alpha_2^{h+1} &\ldots & \alpha^{h+1}_{n}\\
  \vdots &\vdots &\ddots&\vdots \\
  \alpha_1^{k-1} &\alpha_2^{k-1} &\ldots & \alpha^{k-1}_{n}\end{array}\right)$  } \end{equation} is the generator matrix of the twisted Reed-Solomon code $\mathcal C_k(\boldsymbol{\alpha}, t,h,\eta)$.


Note that in general, the twisted Reed-Solomon codes are not MDS. Beelen {\em et al.} got some results on the twisted Reed-Solomon codes as follows:
 \begin{lemma}\label{lem2} \cite[Theorem 17]{BPR} {\rm  Let $\Bbb F_s \subset \Bbb F_q$ be a proper subfield and $\alpha_{1},\ldots,\alpha_{n}\in \Bbb F_s$. If $\eta \in\Bbb F_q \backslash \Bbb F_s $, then the twisted Reed-Solomon code $\mathcal C_k(\boldsymbol{\alpha}, t,h,\eta)$ is MDS.}
 \end{lemma}

  \begin{lemma} \cite[Theorem 18]{BPR} {\rm  Let  $\alpha_{1},\ldots,\alpha_{n}\in \Bbb F_q$ and $2<k<n-2$. Furthermore, let $H\subseteq \Bbb F_q$ satisfy that the  twisted Reed-Solomon code $\mathcal C_k(\boldsymbol{\alpha}, t,h,\eta)$ is MDS for every $\eta\in H$. Then there are at most $6$ choices of $\eta\in H$ such that $\mathcal C_k(\boldsymbol{\alpha}, t,h,\eta)$ is monomially equivalent to a Reed-Solomon code.

   }
 \end{lemma}

  \begin{lemma} \label{lem3} \cite[Corollary 20]{BPR} {\rm Let $\Bbb F_s \subset \Bbb F_q$ with $|\Bbb F_q \backslash \Bbb F_s|>6$. Let $2<k<n-2$ and $n\le s$. Then there exists $\eta\in \Bbb F_q \backslash \Bbb F_s$ such that $\mathcal C_k(\boldsymbol{\alpha}, t,h,\eta)$ is MDS but not monomially equivalent to a  Reed-Solomon code. }
 \end{lemma}

Throughout the paper, if a code is not monomially equivalent to a Reed-Solomon code, then we call it a code of {\em non-Reed-Solomon type} or a {\em  non-Reed-Solomon code}.

\section{ Twisted Reed-Solomon codes with one-dimensional hull}

Let $\gamma$ be a primitive element of $\Bbb F_q$ and $k\mid (q-1)$. Then $\gamma^{\frac{q-1}{k}}$ generates a subgroup of $\Bbb F_q^*$ of order $k$. Let $\alpha_i=\gamma^{\frac{q-1}{k}i}$ for $1\le i\le k$. One can easily check that  \begin{equation}\label{eq3}
\theta_f=\alpha_1^f+\cdots+\alpha_{k}^f=\left\{
\begin{array}{ll}
  k   &      \mbox{if}\ f\equiv0\pmod {k},\\
0 & \mbox{otherwise}.
\end{array} \right.
\end{equation}


\begin{lemma}\label{lem7}{\rm  Let $q$ be a power of two. If $k$ is a positive integer with $k\mid (q-1)$, $k<(q-1)$, and $h>0$, then there exists a $[2k,k]_q$  twisted Reed-Solomon code $\mathcal C_k(\boldsymbol{\alpha}, t,h,\eta)$ over $\Bbb F_q$ with one-dimensional hull   for
$\boldsymbol{\alpha}=(\alpha_1,\ldots,\alpha_k,\gamma\alpha_1,\ldots,\gamma\alpha_k)$,
where $\gamma$ is a primitive element of $\Bbb F_q$ and $\alpha_i=\gamma^{\frac{q-1}{k}i}$ for $1\le i\le k$.

}


\end{lemma}

{\bf Proof} By Definition \ref{def2}, to make sure that $\mathcal C_k(\boldsymbol{\alpha}, t,h,\eta)$ is a twisted Reed-Solomon code, we need $k\neq q-1$.
From  (1), we recall that $G$ is a generator matrix of  the twisted Reed-Solomon code $\mathcal C_k(\boldsymbol{\alpha}, t,h,\eta)$ over $\Bbb F_q$.


Let
\begin{eqnarray*}\resizebox{8.5cm}{!}{$A_{\beta}=\left( \begin{array}{cccccc}
1 & 1& \ldots  & 1 &1\\
\beta\alpha_1& \beta\alpha_2& \ldots   &  \beta\alpha_{k-1}  & \beta\alpha_k \\
\vdots& \vdots& \ldots  & \vdots &\vdots\\
(\beta\alpha_1)^{k-1}& (\beta\alpha_2)^{k-1} &\cdots  & (\beta\alpha_{k-1})^{k-1} &(\beta\alpha_k)^{k-1} \\
\end{array} \right).$}
\end{eqnarray*}
By  (\ref{eq3}), we have
\begin{eqnarray*} \label{eq5}A_{\beta}A_{\beta}^T=\left( \begin{array}{ccccccc}
k &0&0& \ldots  & 0 &0\\
0& 0&0& \ldots   & 0  & \beta^k k \\
\vdots&\vdots& \vdots& \ldots  & \vdots &\vdots\\
0&0&\beta^k k&\cdots  & 0&0 \\
0&\beta^k k&0&\cdots  & 0&0 \\
\end{array} \right).
\end{eqnarray*}
Let $C_{\beta}=A_{\beta}+B_{\beta}$, where
\begin{eqnarray}\resizebox{8.9cm}{!}{$
B_{\beta}=\left( \begin{array}{cccccc}
0 & 0& \ldots  & 0 &0\\
\vdots& \vdots& \ldots  & \vdots &\vdots\\
\eta(\beta\alpha_1)^{k-1+t}& \eta(\beta\alpha_2)^{k-1+t}& \ldots   &  \eta(\beta\alpha_{k-1})^{k-1+t}  &\eta(\beta\alpha_k)^{k-1+t} \\
\vdots& \vdots& \ldots  & \vdots &\vdots\\
0 & 0& \ldots  & 0 &0\\
\end{array} \right)\begin{array}{l} \\   \\\leftarrow(h+1)th\\  \\ \\  \end{array}$}.\end{eqnarray}
Let $\theta_j=\sum_{i=1}^{n}\alpha_i^j$ and $l=k-1+t$, then we have \\
\begin{eqnarray*}&&
C_{\beta}C_{\beta}^T
= \begin{pmatrix}
k &0& \ldots  & 0 &0\\
0& 0& \ldots   & 0  & \beta^k k \\
\vdots& \vdots& \ldots  & \vdots &\vdots\\
0&\beta^k k&\cdots  & 0&0 \\
\end{pmatrix} \\
&+&
\resizebox{9cm}{!}{$\begin{pmatrix}
0 & 0& \ldots  &0& \eta\beta^l\theta_{l}&0  &\ldots&0\\
0 & 0& \ldots  &0& \eta\beta^{l+1}\theta_{l+1}&0 &\ldots&0\\
\vdots & \vdots& \ldots  &\vdots& \vdots & \vdots& \ldots  &\vdots\\
0 & 0& \ldots  &0& \eta\beta^{l+h-1}\theta_{l+h-1}&0  &\ldots& 0\\
\eta\beta^l\theta_{l}& \eta\beta^{l+1}\theta_{l+1}& \ldots  &\eta\beta^{l+h-1}\theta_{l+h-1}&
2\eta\beta^{l+h}\theta_{l+h}+\eta^2\beta^{2l}\theta_{2l}&\eta\beta^{l+h+1}\theta_{l+h+1}&\ldots&\eta\beta^{l+k-1}\theta_{l+k-1}\\
0 &0& \ldots  &0& \eta\beta^{l+h+1}\theta_{l+h+1}&0 &\ldots&0\\
\vdots & \vdots& \ldots  &\vdots & \vdots & \vdots& \ldots  &\vdots\\
0 & 0& \ldots  &0 & \eta\beta^{l+k-1}\theta_{l+k-1}&0 &\ldots&0
\end{pmatrix}$.}
\end{eqnarray*} Since every $\theta_t$ for $l\leq t\leq l+k-1$ is zero except exactly one $\theta_{t'}$, we
can rewrite 
\begin{eqnarray*}&&C_{\beta}C_{\beta}^T=\left( \begin{array}{cccccc}
k &0& \ldots  & 0 &0\\
0& 0& \ldots   & 0  & \beta^k k \\
\vdots& \vdots& \ldots  & \vdots &\vdots\\
0&\beta^k k&\cdots  & 0&0 \\
\end{array} \right)\\
&+&\left( \begin{array}{cccccccc}
0 &\ldots&0& \ldots  & 0 &\ldots&0\\
\vdots&& \vdots& \ldots  & \vdots &&\vdots\\
0&\ldots &0& \ldots   & *_{\beta}  & \ldots&0 \\
\vdots&& \vdots& \ldots  & \vdots &&\vdots\\
0&\ldots &*_{\beta}& \ldots   & \Delta_{\beta}  &\ldots& 0 \\
\vdots&& \vdots& \ldots  & \vdots &&\vdots\\
0&\ldots&0&\cdots  & 0&\ldots&0 \\
\end{array} \right),\end{eqnarray*}
where $\ast_{\beta}$ and $\Delta_{\beta}$ are all elements in $\Bbb F_q$, the  $\ast_{\beta}$ and $\Delta_{\beta}$ are respectively entries located in the $(i+1,h+1)th$, $(h+1,i+1)th$ and $(h+1,h+1)th$ positions, and the other elements are all zero.

Let $G=[C_1: C_{\gamma}]$ and $h>0$. Then \begin{eqnarray*}&&GG^T=C_1C_1^T+C_{\gamma}C_{\gamma}^T
\\
&=&\left( \begin{array}{cccccc}
2k &0& \ldots  & 0 &0\\
0& 0& \ldots   & 0  & (1+\gamma^k) k \\
\vdots& \vdots& \ldots  & \vdots &\vdots\\
0&(1+\gamma^k) k&\cdots  & 0&0 \\
\end{array} \right)\\
&+&\left( \begin{array}{cccccccc}
0 &\ldots&0& \ldots  & 0 &\ldots&0\\
\vdots&& \vdots& \ldots  & \vdots &&\vdots\\
0&\ldots &0& \ldots   & *_{1} + *_{\gamma}& \ldots&0 \\
\vdots&& \vdots& \ldots  & \vdots &&\vdots\\
0&\ldots &*_{1} + *_{\gamma}& \ldots   &  \Delta_1+\Delta_{\gamma}  &\ldots& 0 \\
\vdots&& \vdots& \ldots  & \vdots &&\vdots\\
0&\ldots&0&\cdots  & 0&\ldots&0 \\
\end{array} \right).\end{eqnarray*}
It is easy to find an elementary matrix $P$ such that \begin{eqnarray*}&&PGG^TP^T =
\left( \begin{array}{cccccc}
2k &0& \ldots  & 0 &0\\
0& 0& \ldots   & 0  & (1+\gamma^k) k \\
\vdots& \vdots& \ldots  & \vdots &\vdots\\
0&(1+\gamma^k) k&\cdots  & 0&0 \\
\end{array} \right)\\
&+&\left( \begin{array}{cccccccc}
0 &\ldots&0& \ldots  & 0 &\ldots&0\\
\vdots&& \vdots& \ldots  & \vdots &&\vdots\\
0&\ldots &0& \ldots   & 0& \ldots&0 \\
\vdots&& \vdots& \ldots  & \vdots &&\vdots\\
0&\ldots &0& \ldots   &  \Delta_1+\Delta_{\gamma}  &\ldots& 0 \\
\vdots&& \vdots& \ldots  & \vdots &&\vdots\\
0&\ldots&0&\cdots  & 0&\ldots&0 \\
\end{array} \right).\end{eqnarray*}

By the given conditions,  we have $k\mid (q-1)$, $k< (q-1)$, $q$ is even, and $\gamma $ is a primitive element of $\Bbb F_q$. Hence, $\gamma ^k+1\neq0$ and $2k=0$.  Then we have  $\mbox{rank}(GG^T)=k-1$.  The result follows from Lemma 2.1.
$\blacksquare$


\begin{lemma}\label{lem7}{\rm  Let $\Bbb F_q$ be a finite field of odd order $q$ and $k$ be  a positive integer with $k\mid (q-1)$ and $2<k<{(q-1)}/{2}$. If  $h>1$, then there exists a $[2k,k-1]_q$  twisted Reed-Solomon code $\mathcal C_{k-1}(\boldsymbol{\alpha}, t,h,\eta)$ over $\Bbb F_q$ with one-dimensional hull   for
$\boldsymbol{\alpha}=(\alpha_1,\ldots,\alpha_k,\gamma\alpha_1,\ldots,\gamma\alpha_k)$,
where $\gamma$ is a primitive element of $\Bbb F_q$ and $\alpha_i=\gamma^{\frac{q-1}{k}i}$ for $1\le i\le k$.

}

\end{lemma} 

{\bf Proof} Let \begin{equation}\resizebox{8cm}{!}{$D_{\beta}=\left( \begin{array}{cccccc}
1 & 1& \ldots  & 1 &1\\
\beta\alpha_1& \beta\alpha_2& \ldots   &  \beta\alpha_{k-1}  & \beta\alpha_k \\
\vdots& \vdots& \ldots  & \vdots &\vdots\\
(\beta\alpha_1)^{k-2}& (\beta\alpha_2)^{k-2} &\cdots  & (\beta\alpha_{k-1})^{k-2} &(\beta\alpha_k)^{k-2} \\
\end{array} \right)$.}\end{equation} Then \begin{eqnarray*}&&
D_{\beta}D_{\beta}^T
= \begin{pmatrix}
k &0 &0& \ldots  & 0 &0\\
0 &0 &0& \ldots  & 0 &0\\
0 &0& 0& \ldots   & 0  & \beta^k k \\
\vdots& \vdots& \ldots  & \vdots &\vdots\\
0 &0&\beta^k k&\cdots  & 0&0 \\
\end{pmatrix}. \end{eqnarray*}
Let $H_{\beta}=D_{\beta}+E_{\beta}$, where
\begin{eqnarray*}\resizebox{8.9cm}{!}{$
E_{\beta}=\left( \begin{array}{cccccc}
0 & 0& \ldots  & 0 &0\\
\vdots& \vdots& \ldots  & \vdots &\vdots\\
\eta(\beta\alpha_1)^{k-2+t}& \eta(\beta\alpha_2)^{k-2+t}& \ldots   &  \eta(\beta\alpha_{k-1})^{k-2+t}  &\eta(\beta\alpha_k)^{k-2+t} \\
\vdots& \vdots& \ldots  & \vdots &\vdots\\
0 & 0& \ldots  & 0 &0\\
\end{array} \right)\begin{array}{l} \\   \\\leftarrow(h+1)th\\  \\ \\  \end{array}$}.\end{eqnarray*}

 Let $G=[H_1:H_{\gamma}]$. By the proof of Lemma 3.1,  \begin{eqnarray*}&&GG^T=H_1H_1^T+H_{\gamma}H_{\gamma}^T
\\
&=&\left( \begin{array}{cccccccc}
2k &0 &0& \ldots  & 0 &0&0\\
0 &0 &0& \ldots  & 0 &0&0\\
0 &0& 0& \ldots   & 0  & (1+\gamma^k) k &0\\
\vdots& \vdots& \ldots  & \vdots &\vdots\\
0 &0&(1+\gamma^k) k&\cdots  & 0&0&0 \\
\end{array} \right)\\
&+&\left( \begin{array}{cccccccc}
0 &\ldots&0& \ldots  & 0 &\ldots&0\\
\vdots&& \vdots& \ldots  & \vdots &&\vdots\\
0&\ldots &0& \ldots   & *_{1} + *_{\gamma}& \ldots&0 \\
\vdots&& \vdots& \ldots  & \vdots &&\vdots\\
0&\ldots &*_{1} + *_{\gamma}& \ldots   &  \Delta_1+\Delta_{\gamma}  &\ldots& 0 \\
\vdots&& \vdots& \ldots  & \vdots &&\vdots\\
0&\ldots&0&\cdots  & 0&\ldots&0 \\
\end{array} \right).\end{eqnarray*} Note that $q$ is odd and $k<\frac{q-1}{2}$. We have $2k\neq 0$ and $1+\gamma^k\neq0$. By the same process of the proof of Lemma 3.1, the rank of $GG^T$ is $k-1$ and result follows from Lemma 2.1.
$\blacksquare$

\begin{remark} By the process of Lemma 3.2, we can also construct some twisted Reed-Solomon codes with small hulls.
\end{remark}

An effective method for construction of twisted Reed-Solomon codes with MDS property
is to use the lifting  of the finite field (refer to \cite{BPR}).   Hence, we obtain the following theorem by Lemmas 2.5, 2.7, 3.1 and 3.2.

\begin{theorem} {\rm Let $q$ be a power of a prime and $\Bbb F_s \subset \Bbb F_q$ with
$|\Bbb F_q \backslash \Bbb F_s|>6$. Suppose that  $k$ is a positive integer with $k\mid (q-1)$. 


(1) If $q$ is even and  $2<k< (s-1)$,
then there exists a $[2k, k]_q$  MDS non-Reed-Solomon code with one-dimensional hull. 

(2) If $q$ is odd and  $2<k< (s-1)/2$,
then there exists a $[2k, k-1]_q$  MDS non-Reed-Solomon code with one-dimensional hull.

}
\end{theorem}

%
%


In the following, we will present some examples to show our main results.
\begin{example}
{\rm
(1) Let $q=2^4=16$, $k=5$, and $\gamma$ be a primitive element of $\Bbb F_{q}$. Consider a twisted Reed-Solomon code $\mathcal C_5(\boldsymbol{\alpha}, 1,3,\eta)$, when $\boldsymbol{\alpha}=(1,\gamma^{3}, \gamma^{6}, \gamma^{9},  \gamma^{12},\gamma,\gamma\gamma^{3}, \gamma\gamma^{6}, \gamma\gamma^{9},\gamma\gamma^{12})$ and $\eta=\gamma^i\in \Bbb F_{16}$. 
By Lemma 3.1, $\mathcal C_5(\boldsymbol{\alpha}, 1,3,\gamma^i)$ has one-dimensional hull for all $i$. By Magma, it follows that
the codes $\mathcal C_5(\boldsymbol{\alpha}, 1,3,\eta)$ are not MDS codes, which have parameters $[10,5,5]_{16}$.

(2) Let $q=2^8=256$, $k=5$, and $w$ be a primitive element of $\Bbb F_{q}$ and $\gamma=w^{17}\in \Bbb F_{16}$. Consider a twisted Reed-Solomon code $\mathcal C_5(\boldsymbol{\alpha}, 1,3,\eta)$, when $\boldsymbol{\alpha}=(1,\gamma^{3}, \gamma^{6}, \gamma^{9},  \gamma^{12},\gamma,\gamma\gamma^{3}, \gamma\gamma^{6}, \gamma\gamma^{9},\gamma\gamma^{12})$ and $\eta=w^i\in \Bbb F_{256}$. 
By Lemma 3.1, $\mathcal C_5(\boldsymbol{\alpha}, 1,3,w^i)$ has one-dimensional hull for all $i$. By Magma and Lemmas 2.5 and 2.7, there exists an integer $i$ with $17\nmid i$ such that 
the code $\mathcal C_5(\boldsymbol{\alpha}, 1,3,\eta)$ is an   MDS  non-Reed-Solomon code with parameters $[10,5,6]_{256}$.

}

\end{example}

\begin{example}
{\rm
(1) Let $q=3^4=81$, $k=5$, and $\gamma$ be a primitive element of $\Bbb F_{q}$. Consider a twisted Reed-Solomon code $\mathcal C_4(\boldsymbol{\alpha}, 2,2,\eta)$, when $\boldsymbol{\alpha}=(1,w, w^{2}, w^{3},  w^{4},\gamma,\gamma w, \gamma w^{2}, \gamma w^{3},\gamma w^{4})$ and $\eta=\gamma^i\in \Bbb F_{81}$ and $w=\gamma^{16}$. 
By Lemma 3.2, $\mathcal C_4(\boldsymbol{\alpha}, 2,2,\gamma^i)$ has one-dimensional hull for all $i$. By Magma, it follows that
the codes $\mathcal C_4(\boldsymbol{\alpha}, 2,2,\gamma^i)$ are MDS with parameters $[10,4]_{81}$ when $\eta$ belongs to $H=\{\gamma^j: j=0,6,16,22,32,38,48,54,64,70\}$.
  Since $|H|>6$, by Lemma 2.6, there exists $\eta\in H$ such that $\mathcal C_4(\boldsymbol{\alpha}, 2,2,\eta)$ is a non-Reed-Solomon code. As a result, there exists a $[10,4]_{81}$   MDS non-Reed-Solomon code with one-dimensional hull.

(2) Let $q=3^8=6561$, $k=5$, and $\theta$ be a primitive element of $\Bbb F_{q}$, $\gamma=\theta^{82}\in \Bbb F_{81}$, and $w=\gamma^{16}$. Consider a twisted Reed-Solomon code $\mathcal C_4(\boldsymbol{\alpha}, 2,2,\eta)$, when $\boldsymbol{\alpha}=(1,w, w^{2}, w^{3},  w^{4},\gamma,\gamma w, \gamma w^{2}, \gamma w^{3},\gamma w^{4})$ and $\eta=\theta^i\in \Bbb F_{6561}$. 
By Lemma 3.2, $\mathcal C_4(\boldsymbol{\alpha}, 2,2,\gamma^i)$ has one-dimensional hull for all $i$. By Magma and Lemmas 2.5 and 2.7, there exists an integer $i$ with $82\nmid i$ such that 
the code $\mathcal C_4(\boldsymbol{\alpha}, 2,2,\gamma^i)$ is an  MDS  non-Reed-Solomon code with parameters $[10,4,7]_{6561}$.

}

\end{example}

\section{Concluding remarks}

For a given linear code, in general case it is hard to show that  if the code is monomially equivalent to a Reed-Solomon code with the same parameters.
In this paper, we applied twisted Reed-Solomon codes to construct  some MDS codes, which  have one-dimensional hull and are not monomially equivalent to Reed-Solomon codes. We also presented some examples by using Magma.





%

\section*{Acknowledgment}


The author is very grateful to the reviewers and the Associate Editor for their valuable comments and suggestions to improve the quality of this paper.

\ifCLASSOPTIONcaptionsoff
  \newpage
\fi

\end{document}